# A Practical Review of Mechanistic Interpretability for Transformer-Based Language Models

### v2

@@ 7.1 Feature Study
Most MI studies rely on the linear representation hypothesis Mikolov et al. (2013); Pennington et al. (2014) for performing feature studies. The hypothesis states that features are encoded linearly in the representation space of the model. For example, SAEs adopt this hypothesis by assuming that activations can be expressed as sparse linear combinations of feature vectors, while probing assumes that a linear model can detect the presence or absence of a feature in activations. Specifically, the linear representation hypothesis posits that neural networks have two properties: linearity, i.e. the network’s activation space consists of meaningful (linear) vectors, each representing a feature, and decomposability, i.e., network activations can be decomposed and described in terms of these independent features. Notably, this hypothesis predates MI and has been extensively studied to understand how high-level concepts are stored in word embeddings Mikolov et al. (2013), latent representations of variational autoencoders Wang et al. (2024d), and LM representations Dai et al. (2022); Radford et al. (2017). A well-known example of the hypothesis is the presence of semantic relationships such as $\overrightarrow{man\vphantom{k}}-\overrightarrow{woman\vphantom{k}}\approx\overrightarrow{king}-\overrightarrow{queen\vphantom{k}}$ within word embeddings, providing evidence that high-level concepts (e.g., gender) are represented as specific directions in the activation space. An increasing body of work has found evidence supporting the hypothesis by discovering features encoded as directions within LM representations such as sentiment Tigges et al. (2023), space and time Gurnee & Tegmark (2023), language Bricken et al. (2023), truth Marks & Tegmark (2023), and refusal Arditi et al. (2024). However, some studies have also found the existence of non-linear features that contradict the linear representation hypothesis. For instance, Engels et al. (2024) discovered multi-dimensional features, such as circular representations for days of the week and months of the year, demonstrating that some features are inherently irreducible to a single dimension. This finding contradicts the previous notion of linear representation hypothesis where each feature is represented by one-dimensional directions. Subsequently, Olah (2023) pointed out that since multi-dimensional features remain mathematically linear, existing techniques should be capable of detecting them with a slight change. However, Csordás et al. (2024) found that gated recurrent neural networks (RNNs) encode sequential information using non-linear “onion representations”, where tokens are represented by magnitude rather than direction, leading to layered features that do not reside in distinct linear subspaces. Given these contradictory findings, further research is crucial to either validate or refute the linear representation hypothesis, as it forms the foundation of much feature analysis.
@@ 7.2.1 Interpreting LM Behaviors
Circuit studies have identified a range of behaviors across LMs of different sizes, as listed in Table 3. These circuits are defined at varying levels of granularity, where nodes in the computation graph are defined as the outputs of MHA and FF sublayers (Olsson et al., 2022; Wang et al., 2022a) or as SAE features Marks et al. (2024b); Cunningham et al. (2023); Kissane et al. (2024). In addition, Merullo et al. (2023) and Quirke et al. (2024) showed that the same components (e.g., induction heads) are reused by different circuits (e.g., IOI and induction circuits) to implement different tasks, demonstrating the generalizability of interpreted components. Beyond circuit discovery, some studies Heimersheim & Janiak (2023); Marks et al. (2024b) have shown that the identified circuits can be enhanced for better performance, highlighting the potential for practical downstream applications. For instance, Marks et al. (2024b) proposed a technique, Spurious Human-interpretable Feature Trimming (SHIFT), which improves the generalization of a classifier by removing the spurious features in the discovered circuit. This approach paves the way for new possibilities in human-AI collaboration.
@@
Subject-verb Agreement Marks et al. (2024b) Completes sentences like “The keys in the cabinet” with “are” as opposed to “is” $<100$ SAE features Pythia-70M
Arithmetic Calculation Nikankin et al. (2024) $51+4\xrightarrow{}$ 5; $48-4\xrightarrow{}$ 4; $82*4\xrightarrow{}$ 8; $28/4\xrightarrow{}$ 2 6, 6, 20, and 6 attention heads, along with all FF sub-layers, correspond to Addition, Subtraction, Multiplication, and Division, respectively Llama3-8B
@@ 7.2.2 Interpreting LM components
The LayerNorms Ba (2016) are primarily employed to stabilize and accelerate the training of LMs. However, their role in model computation during inference remains less understood. Many MI studies assume that LayerNorm layers do not play a significant role during inference and thus often ignore their computation. To investigate this assumption, Heimersheim (2024) demonstrated that the LayerNorm in a pre-trained GPT2-small can be removed by fine-tuning the model on a small fraction of the training data without the LayerNorm, suggesting that it may not be essential for the core computational process. In contrast, other studies have revealed that LayerNorm can have additional functionalities; for instance, Winsor (2020) showed that LayerNorm can serve as a general-purpose non-linear function capable of performing various classification tasks, while Stolfo et al. (2024) proposed that it is instrumental in implementing confidence regularization within LMs. These findings highlight that while LayerNorm is critical during training, its exact contributions during inference merit further exploration.
@@ 7.5 Findings on Learning Dynamics
Phase Changes during LM Training
@@
Effects of Post-training
@@ Acknowledgements
DR and ZY were sponsored by the National Science Foundation (#2311468/#2423813) and the Department of Computer Science at the College of Computing and Engineering at George Mason University. We appreciate comments from readers of the earlier version of this paper.

### v4

@@ 8.1 Feature Study
Most MI studies rely on the linear representation hypothesis Mikolov et al. (2013); Pennington et al. (2014) for performing feature studies. The hypothesis states that features are encoded linearly in the representation space of the model. For example, SAEs adopt this hypothesis by assuming that activations can be expressed as sparse linear combinations of feature vectors, while probing assumes that a linear model can detect the presence or absence of a feature in activations. Specifically, the linear representation hypothesis posits that neural networks have two properties: linearity, i.e., the network’s activation space consists of meaningful (linear) vectors, each representing a feature, and decomposability, i.e., network activations can be decomposed and described in terms of these independent features. Notably, this hypothesis predates MI and has been extensively studied to understand how high-level concepts are stored in word embeddings Mikolov et al. (2013), latent representations of variational autoencoders Wang et al. (2024d), and LM representations Dai et al. (2022); Radford et al. (2017). A well-known example of the hypothesis is the presence of semantic relationships such as $\overrightarrow{man\vphantom{k}}-\overrightarrow{woman\vphantom{k}}\approx\overrightarrow{king}-\overrightarrow{queen\vphantom{k}}$ within word embeddings, providing evidence that high-level concepts (e.g., gender) are represented as specific directions in the activation space. An increasing body of work has found evidence supporting the hypothesis by discovering features encoded as directions within LM representations, such as sentiment Tigges et al. (2023), space and time Gurnee & Tegmark (2023), language Bricken et al. (2023), truth Marks & Tegmark (2023), and refusal Arditi et al. (2024). However, some studies have also found the existence of non-linear features that contradict the linear representation hypothesis. For instance, Engels et al. (2024) discovered multi-dimensional features, such as circular representations for days of the week and months of the year, demonstrating that some features are inherently irreducible to a single dimension. This finding contradicts the previous notion of linear representation hypothesis where each feature is represented by one-dimensional directions. Subsequently, Olah (2023) pointed out that since multi-dimensional features remain mathematically linear, existing techniques should be capable of detecting them with a slight change. However, Csordás et al. (2024) found that gated recurrent neural networks (RNNs) encode sequential information using non-linear “onion representations”, where tokens are represented by magnitude rather than direction, leading to layered features that do not reside in distinct linear subspaces. Given these contradictory findings, further research is crucial to either validate or refute the linear representation hypothesis, as it forms the foundation of much feature analysis.
@@ 8.2.1 Interpreting LM Behaviors
Circuit studies have identified a range of behaviors across LMs of different sizes, as listed in Table 8.2.1. These circuits are defined at varying levels of granularity, where nodes in the computation graph are defined as the outputs of MHA and FF sublayers (Olsson et al., 2022; Wang et al., 2022a) or as SAE features Marks et al. (2024b); Cunningham et al. (2023); Kissane et al. (2024). In addition, Merullo et al. (2023) and Quirke et al. (2024) showed that the same components (e.g., induction heads) are reused by different circuits (e.g., IOI and induction circuits) to implement different tasks, demonstrating the generalizability of interpreted components. Beyond circuit discovery, some studies Heimersheim & Janiak (2023); Marks et al. (2024b) have shown that the identified circuits can be enhanced for better performance, highlighting the potential for practical downstream applications. For instance, Marks et al. (2024b) proposed a technique, Spurious Human-interpretable Feature Trimming (SHIFT), which improves the generalization of a classifier by removing the spurious features in the discovered circuit. This approach paves the way for new possibilities in human-AI collaboration.
@@
Subject-verb Agreement Marks et al. (2024b) Completes sentences like “The keys in the cabinet” with “are” as opposed to “is” <100<100 SAE features Pythia-70M
Arithmetic Calculation Nikankin et al. (2024) $1+4\xrightarrow{}5$; $8-4\xrightarrow{}4$; $2*4\xrightarrow{}8$; $8/4\xrightarrow{}2$ 6, 6, 20, and 6 attention heads, along with all FF sub-layers, correspond to Addition, Subtraction, Multiplication, and Division, respectively Llama3-8B
@@ 8.2.2 Interpreting LM components
The LayerNorms Ba et al. (2016) are primarily employed to stabilize and accelerate the training of LMs. However, their role in model computation during inference remains less understood. Many MI studies assume that LayerNorm layers do not play a significant role during inference and thus often ignore their computation. To investigate this assumption, Heimersheim (2024) demonstrated that the LayerNorm in a pre-trained GPT2-small can be removed by fine-tuning the model on a small fraction of the training data without the LayerNorm, suggesting that it may not be essential for the core computational process. In contrast, other studies have revealed that LayerNorm can have additional functionalities; for instance, Winsor (2020) showed that LayerNorm can serve as a general-purpose non-linear function capable of performing various classification tasks, while Stolfo et al. (2024) proposed that it is instrumental in implementing confidence regularization within LMs. These findings highlight that while LayerNorm is critical during training, its exact contributions during inference merit further exploration.
@@ 8.5 Findings on Learning Dynamics
8.5.1 Phase Changes during LM Training
@@
8.5.2 Effects of Post-training
@@ 8.6.3 Others
8.7 Benchmarks for Evaluating Interpretability Techniques
8.7.1 Benchmarks for Feature Study Techniques
For feature study, several benchmarks have been developed to evaluate how well various techniques (discussed in Section 5) can identify meaningful feature vectors for specific concepts, i.e., targeted feature study. For instance, CausalGym (Arora et al., 2024) introduces a suite of linguistic tasks that test the effectiveness of the technique in discovering the feature vector corresponding to a given concept, and shows that DAS outperforms alternative approaches such as probing. Similarly, Resolving Attribute–Value Entanglements in Language Models (RAVEL) (Huang et al., 2024a) benchmarks targeted feature study methods on their ability to localize and disentangle specific attributes of entities (e.g., “Paris is on the continent of”). Their results again show that DAS achieves the strongest performance, outperforming methods like SAE and probing. Mechanistic Interpretability Benchmark (MIB) (Mueller et al., 2025) is another benchmark that provides a broader benchmark covering both targeted and open-ended feature study. Their benchmark also finds DAS to be the best technique for isolating the feature vector for a given concept. For open-ended feature study, however, MIB reports a surprising finding: SAE features are no more effective than individual neurons at isolating meaningful features, contrary to prior claims. SAEBench (Karvonen et al., 2025) is another benchmark that provides a comprehensive evaluation suite for comparing various SAE architectures and training setups. Specifically, it proposes to evaluate SAEs across eight diverse metrics, including interpretability, feature disentanglement, and practical applications such as unlearning. Their benchmark results indicate that Matryoshka SAEs achieve the best overall performance, particularly in feature disentanglement. Finally, Function INterpretation and Description (FIND) Schwettmann et al. (2023) is a benchmark that focuses on the Explain step of open-ended feature study, evaluating the correctness of interpretability LM agents in describing the latent functions implemented by model components.
8.7.2 Benchmarks for Circuit Study Techniques
Several benchmarks have also been proposed to evaluate the effectiveness of techniques for the localization step of circuit study. Some works employ collections of synthetic (Lindner et al., 2024) or semi-synthetic (Gupta et al., 2024) transformers with known circuits, where the availability of ground truth makes evaluation more straightforward. Results from these studies show that ACDC and EAP-IG achieve the strongest performance. However, concerns remain that such synthetic benchmarks may not faithfully capture the behavior of standard pre-trained transformers. To address this, MIB (Mueller et al., 2025) evaluates circuit localization techniques on both semi-synthetic and standard pre-trained transformer models, where they also found that EAP-IG outperforms other approaches like edge patching.
@@ Acknowledgements
DR and ZY were sponsored by the National Science Foundation (#2311468/#2423813) and the Department of Computer Science at the College of Computing and Engineering at George Mason University. We appreciate comments from readers of the earlier version of this paper, and the clarification from Johnny Lin on the usage of Neuronpedia.